# Basis Selection: Low-Rank Decomposition of Pretrained Large Language Models for Target Applications

### v1

@@ 12       Tune the learnable parameters;
22
23for each layer do
24       Compute the low rank matrix $\widetilde{W}$ based on equation (6);
25       $[U^{\prime},S^{\prime},V^{\prime}]=\text{SVD}(\widetilde{W})$;
@@ 33       FineTune the new model $M^{\prime}$;
Algorithm 1 Basis Selection Algorithm
@@ 4 Experiments
Figure 3: Pass@1 accuracy and model size of Llama 2-7B compressed by various algorithms for the mathematical reasoning task.
@@
Figure 4: Pass@1 accuracy and model size of Llama 2-13B compressed by various algorithms for the mathematical reasoning task.
4.1 Evaluation Methodology
The performance of low-rank compression algorithms is evaluated on two tasks: mathematical reasoning and code generation. For each task, Llama 2-7B and Llama 2-13B models [17] are first finetuned on a training dataset and then compressed using a compression algorithm. The compressed models are further finetuned before evaluation.
For the mathematical reasoning task, we utilize two evaluation datasets: GSM8K [6] and Hendrycks’ MATH [10]. The GSM8K dataset comprises verbally described mathematical questions, containing 1,319 samples used for evaluation. The Hendrycks’ MATH dataset covers more complex topics such as linear algebra and geometry, consisting of 5,000 question-answer pairs used for evaluation. Due to its complexity, the Hendrycks’ MATH dataset necessitates more sophisticated computations and reasoning, resulting in lower accuracy compared to GSM8K.
For the code generation task, we use two evaluation datasets: MBPP [2] and HumanEval [3]. Both datasets evaluate the models’ ability to generate Python code. MBPP comprises 500 code generation questions, while HumanEval includes 164 code generation questions.
We compare our proposed Basis Selection with state-of-the-art low-rank compression algorithms, specifically SVD and FWSVD. SVD is widely used in previous model compression studies [16, 1, 14, 18, 13, 7, 20, 15]. FWSVD [11] enhances SVD by evaluating the importance of model weight parameters.
Our work focuses exclusively on low-rank compression. Comparing our approach with model compression algorithms from other areas, such as quantization, sparsity & pruning, and knowledge distillation, is beyond the scope of this study. Ultimately, compression algorithms from different areas can be combined to achieve better compression results.
@@ 4.1 Evaluation Methodology
Figure 5: Pass@1 accuracy and model size of Llama 2-7B compressed by various algorithms for the code generation task.
@@
Figure 6: Pass@1 accuracy and model size of Llama 2-13B compressed by various algorithms for the code generation task.
@@
Figure 7: Ablation study: Effect of varying the additional dimension of Basis Selection on compressing Llama 2-7B for the mathematical reasoning task.
@@
Figure 8: Ablation study: Effect of varying the pruning times of Basis Selection on compressing Llama 2-7B for the mathematical reasoning task.
4.2 Results on Mathematical Reasoning
Figures 3(a) and (b) depict the performance of various compression algorithms on the Llama 2-7B model for the mathematical reasoning task. We evaluate the models’ accuracy (Pass@1) at different compression ratios (original model size vs. compressed model size). For low compression ratios (below 6), all methods achieve similar accuracy. However, our Basis Selection method significantly outperforms SVD and FWSVD at higher compression ratios. For instance, at a 7x compression ratio, Basis Selection achieves around 43% and 10% accuracy on GSM8K and MATH datasets, respectively, while FWSVD drops below 2% accuracy on both datasets and SVD reaches only 27% and 5% accuracy on GSM8K and MATH, respectively. This highlights the effectiveness of Basis Selection for deep compression, especially when aiming for aggressive model size reduction.
Similar trends emerge for the larger Llama 2-13B model in Figures 4(a) and (b). Once again, Basis Selection significantly outperforms SVD and FWSVD at compression ratios exceeding 6. At a 7x compression ratio, Basis Selection achieves 47% and 11% accuracy on GSM8K and MATH datasets, respectively, demonstrating its advantage. This is in stark contrast to SVD’s performance (23% and 5% accuracy on GSM8K and MATH) and FWSVD’s near-complete accuracy drop (around 2% on both datasets). These results solidify Basis Selection’s effectiveness for deep compression across different model sizes.
4.3 Results on Code Generation
Similar results extend to code generation tasks (Figures 5 and  6). For both Llama 2-7B and Llama 2-13B models, all methods perform comparably at lower compression ratios (below 4). However, Basis Selection exhibits clear superiority at higher compression ratios. On Llama 2-7B at a 6x compression ratio, Basis Selection achieves 12% and 8% accuracy on HumanEval and MBPP datasets, respectively, significantly outperforming SVD (5% and 2%) and FWSVD (6% and 2%). Similar trends hold for Llama 2-13B. These findings further solidify Basis Selection’s effectiveness for deep compression across diverse tasks and model sizes.
@@ 4.4 Ablation Study
To analyze the impact of key parameters of Basis Selection, we conducted ablation studies on the additional dimension (denoted by $\tilde{r}$ in equation (6)) and pruning times. The additional dimension compensates for information loss during pruning, especially for deep compression. Figure 7 compares the performance of compressing Llama 2-7B for the mathematical reasoning task with and without an additional dimension of 32. As expected, incorporating an additional dimension improves accuracy at higher compression ratios (above 6). Similarly, pruning the model lightly but multiple times allows it to gradually adapt to the reduction in parameters after each pruning. This is particularly beneficial for achieving extreme compression ratios. Figure 8 demonstrates this effect by comparing the performance of pruning Llama 2-7B 100 times vs. 2 times on the same task. Here, pruning 100 times leads to better accuracy at compression ratios exceeding 10.
@@ 5 Conclusion
The significant size of large language models leads to high inference costs and demands substantial computing resources. To mitigate these issues, we focus on compressing large language models to meet the specific requirements of target applications. Our approach involves examining these models through the lens of matrix factorization. By viewing the weight matrix of large language models as a linear combination of a group of bases, we have identified that pretrained models often contain many redundant bases that are less useful for target applications. To address this, we propose Basis Selection, a compression algorithm that evaluates the importance of each base for target applications and prunes those that are less significant. Experimental results demonstrate that Basis Selection significantly outperforms state-of-the-art low-rank compression algorithms in achieving deep compression. We believe Basis Selection will greatly reduce the inference cost of large language models, making them more accessible and practical for a wider range of applications. This advancement has the potential to democratize the use of large language models, facilitating their adoption and integration across diverse fields and industries.

### v2

@@ 11  Tune the learnable parameters;
21
22for each layer do
23  Compute the low rank matrix $\widetilde{W}$ based on equation equation 3;
[…]
@@ 32  FineTune the new model $M^{\prime}$;
Algorithm 1 Basel Algorithm
In the first term, $\mathbf{u_{i}}\mathbf{v_{i}}^{\mathrm{T}}$ represents the original bases in the pretrained model. To assess their importance, we initialize their weights $\tilde{s}_{i}$ with their original singular values and then retrain these weights (while keeping the bases fixed) on the training set of the target application. The aim is that, after retraining, the bases important for the target application will have larger singular values, whereas those that are useless or detrimental will have zero or very small singular values. This allows us to identify and prune the less useful bases. Relearning the importance of bases for the target application distinguishes our approach from previous methods. Prior approaches either use the singular values in the original model (Xue et al., 2013; Jaderberg et al., 2014; Denton et al., 2014; Zhang et al., 2015; Povey et al., 2018; Chen et al., 2018; Acharya et al., 2019; Noach & Goldberg, 2020; Sharma et al., 2024) or assess the importance of weight parameters, other than the importance of the bases, to prune them (Hsu et al., 2022). They do not relearn the importance of the bases specifically for the target application. From a signal processing perspective, this first term allows us to adjust the weight for each filter, catering to the needs of the target application.
In the second term, the vectors $\tilde{\mathbf{u_{j}}}$ and $\tilde{\mathbf{v_{j}}}$ are learnable and serve two main purposes. First, due to distributional differences between the pretraining data and the target application, some bases essential for the latter may be missing. These vectors are used to learn such missing bases. Second, although each pruned basis may have little effect individually, their collective removal can lead to a noticeable performance drop. The additional vectors help mitigate this loss. The number of learnable vectors, denoted $\tilde{r}$, is called the additional dimension. From a signal processing perspective, the second term introduces new filters that enhance the model’s performance on the target task. While this term adds new parameters, they do not fully contribute to model size growth. This is because SVD is eventually applied to the sum of the first and second terms, reducing the rank if dependencies exist between the newly learned and original, kept bases.
Algorithm 1 presents our proposed approach. Basel operates on a pretrained or fine-tuned model as input. Similar to Hsu et al. (2022); Sharma et al. (2024), Basel iteratively prunes the bases of all linear layers—excluding the embedding layer—including the $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, $\mathbf{W_{V}}$, and $\mathbf{W_{O}}$ matrices in the attention layers, the linear layers in the feedforward layers, and the output layer. The same compression process is applied across these components. After each pruning step, the learnable parameters $\tilde{s}_{i}$, $\tilde{\mathbf{u}}_{j}$, and $\tilde{\mathbf{v}}_{j}$ are finetuned to compensate for performance degradation. Ultimately, a new weight matrix $\widetilde{\mathbf{W}}$ with a smaller rank $r^{\prime}$ is learned. We perform a standard SVD on it, representing it as the product of matrices $\mathbf{U^{\prime}}$, $\mathbf{S^{\prime}}$, and $\mathbf{V^{\prime}}^{\top}$. We then replace the original layer with two new layers: $\mathbf{S^{\prime}V^{\prime}}^{\top}$ becomes the weight matrix of the first new layer, and $\mathbf{U^{\prime}}$ becomes the weight matrix of the second new layer. This reduces the number of parameters from $nm$ to $(n+m)r^{\prime}$. The new model is subsequently further finetuned to enhance its performance on the target application.
@@ 4 Experiments
4.1 Evaluation Methodology
The performance of low-rank compression algorithms is evaluated on two tasks: mathematical reasoning and code generation. For each task, Llama 2-7B and Llama 2-13B models (Touvron et al., 2023b) are first finetuned on a training dataset and then compressed using a compression algorithm. The compressed models are further finetuned before evaluation.
For the mathematical reasoning task, we utilize two evaluation datasets: GSM8K (Cobbe et al., 2021) and Hendrycks’ MATH (Hendrycks et al., 2021). The GSM8K dataset comprises verbally described mathematical questions, containing 1,319 samples used for evaluation. The Hendrycks’ MATH dataset covers more complex topics such as linear algebra and geometry, consisting of 5,000 question-answer pairs used for evaluation. Due to its complexity, the Hendrycks’ MATH dataset necessitates more sophisticated computations and reasoning, resulting in lower accuracy compared to GSM8K.
For the code generation task, we use two evaluation datasets: MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021a). Both datasets evaluate the models’ ability to generate Python code. MBPP comprises 500 code generation questions, while HumanEval includes 164 questions.
We evaluate our proposed Basel method against state-of-the-art low-rank compression techniques, including SVD and FWSVD. SVD has been widely used in prior work on model compression (Sharma et al., 2024; Acharya et al., 2019; Noach & Goldberg, 2020; Xue et al., 2013; Jaderberg et al., 2014; Denton et al., 2014; Zhang et al., 2015; Povey et al., 2018), and in our experiments, it is applied to compress a model that has been fine-tuned on the target task. FWSVD (Hsu et al., 2022) extends SVD by incorporating gradient-based profiling: it estimates weight importance using gradients computed on the fine-tuning dataset of the target task and leverages this information during decomposition. In addition to low-rank methods, we also compare Basel with leading approaches from other compression paradigms, including QLoRA (Dettmers et al., 2023), a quantization-based method, and FLAP (An et al., 2024), a pruning-based method. QLoRA improves performance by training LoRA adapters on the target task’s fine-tuning dataset. FLAP prunes weights based on their magnitudes and the variance of each input activation dimension, profiled on the target task.
4.2 Evaluating Low-Rank Compression Methods for Mathematical Reasoning
Figures 3 (a) and (b) depict the performance of various low-rank compression algorithms on the Llama 2-7B model for the mathematical reasoning task. We evaluate the models’ accuracy (Pass@1) at different compression ratios (original model size vs. compressed model size). For low compression ratios (below 6), all methods achieve similar accuracy. However, our Basel method significantly outperforms SVD and FWSVD at higher compression ratios. For instance, at a 7x compression ratio, Basel achieves around 46% and 12% accuracy on GSM8K and MATH datasets, respectively, while FWSVD drops below 2% accuracy on both datasets and SVD reaches only 27% and 5% accuracy on GSM8K and MATH, respectively. We also find that Basel, at a compression ratio of 16, achieves better accuracy on both GSM8K and MATH compared to FWSVD and SVD at a compression ratio of 6. This suggests that Basel reduces the model size by up to 2.7 times more than the baseline methods while maintaining similar accuracy. This highlights the effectiveness of Basel for deep compression, especially when aiming for aggressive model size reduction.
@@ 4.2 Evaluating Low-Rank Compression Methods for Mathematical Reasoning
Рис. 3: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms on the mathematical reasoning task.
@@
Рис. 4: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms on the mathematical reasoning task.
@@
Рис. 5: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms on the code generation task.
@@
Рис. 6: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms on the code generation task.
Similar trends emerge for the larger Llama 2-13B model in Figures 4 (a) and (b). Once again, Basel significantly outperforms SVD and FWSVD at compression ratios exceeding 6. At a 7x compression ratio, Basel achieves 47% and 11% accuracy on GSM8K and MATH datasets, respectively, demonstrating its advantage. This is in stark contrast to SVD’s performance (23% and 5% accuracy on GSM8K and MATH) and FWSVD’s performance (around 2% on both datasets). These results solidify Basel’s effectiveness for deep compression across different model sizes.
4.3 Evaluating Low-Rank Compression Methods for Code Generation
Similar results extend to code generation tasks (Figures 5 and  6). For both Llama 2-7B and Llama 2-13B models, all methods perform comparably at lower compression ratios (below 4). However, Basel exhibits clear superiority at higher compression ratios. On Llama 2-7B at a 6x compression ratio, Basel achieves 12% and 9% accuracy on HumanEval and MBPP datasets, respectively, significantly outperforming SVD (5% and 2%) and FWSVD (6% and 2%). Similar trends hold for Llama 2-13B. These findings further solidify Basel’s effectiveness for deep compression across diverse tasks and model sizes.
@@ 4.3 Evaluating Low-Rank Compression Methods for Code Generation
Рис. 7: Pass@1 accuracy and model size of Llama 2-7B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the mathematical reasoning task.
@@
Рис. 8: Pass@1 accuracy and model size of Llama 2-13B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the mathematical reasoning task.
4.4 Combining Low-Rank Compression and Quantization for Math Reasoning
Quantization reduces the precision of weight parameters to shrink model size and is a widely used technique for model compression. When using 8-bit quantization or higher, the performance degradation is typically minimal. However, more aggressive quantization (e.g., 4-bit) often leads to a significant drop in accuracy. We apply both 8-bit and 4-bit quantization to Llama 2-7B and Llama 2-13B models on mathematical reasoning.
For Llama 2-7B, the 8-bit quantized model achieves Pass@1 accuracy of 66.0% on GSM8K and 20.3% on MATH, closely matching the unquantized model’s performance (66.4% and 20.6%, respectively, with bf16). A similar trend holds for Llama 2-13B, where 8-bit quantization results in only a negligible accuracy drop. In contrast, 4-bit quantization causes substantial degradation. On Llama 2-7B, accuracy drops from 66.0% to 39.2% on GSM8K and from 20.3% to 8.8% on MATH. Llama 2-13B exhibits similarly significant losses. These results suggest that relying solely on quantization—especially at lower bit-widths—can be detrimental to performance.
To address this, we explore combining quantization with low-rank compression using our Basel method. Figures 7 and 8 illustrate that combining Basel with 8-bit quantization not only achieves a smaller model size but also significantly outperforms 4-bit quantization in accuracy. On compressing Llama 2-7B, Basel + 8-bit achieves 59.6% on GSM8K and 18.1% on MATH with a 3.28 GB model size, compared to 4-bit quantization’s 39.2% and 8.8% with a larger 3.37 GB model. On compressing Llama 2-13B, Basel + 8-bit yields 61.9% on GSM8K and 17.5% on MATH with a 6.13 GB size, while 4-bit quantization achieves only 52.8% and 12.5% with 6.51 GB.
We further compare Basel + 8-bit quantization against two additional approaches: (1) QLoRA with 4-bit quantization and (2) combining 8-bit quantization with FLAP, a state-of-the-art pruning technique. As shown in Figures 7 and 8, Basel + 8-bit consistently outperforms both in terms of accuracy and size. For instance, on compressing Llama 2-7B, it improves accuracy by 5.5% on both GSM8K and MATH compared to QLoRA, while reducing the model size by an additional 0.13 GB.
(a) HumanEval
(b) MBPP
Рис. 9: Pass@1 accuracy and model size of Llama 2-7B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the code generation task.
(a) HumanEval
(b) MBPP
Рис. 10: Pass@1 accuracy and model size of Llama 2-13B compressed using quantization alone, as well as quantization combined with low-rank methods or pruning, on the code generation task.
4.5 Combining Low-Rank Compression and Quantization for Code Generation
We also compare Basel + 8-bit quantization to three alternative methods on the code generation task: (1) standard 4-bit quantization, (2) QLoRA with 4-bit quantization, and (3) FLAP combined with 8-bit quantization. Figures 9 and 10 present the results for compressing Llama 2-7B and Llama 2-13B. On both models, Basel + 8-bit quantization consistently outperforms 4-bit quantization while achieving smaller model sizes. For example, on Llama 2-7B, it improves accuracy by 12.8% on HumanEval and 6.4% on MBPP, with a model size similar to that of the 4-bit version. This further highlights the benefit of combining low-rank compression with quantization.
Basel + 8-bit quantization also significantly outperforms both QLoRA (4-bit) and FLAP + 8-bit quantization on Llama 2-7B and Llama 2-13B. For instance, on compressing Llama 2-7B, it achieves 6.7% higher accuracy on HumanEval and 1.0% higher on MBPP compared to QLoRA, while also reducing the model size by an additional 0.17 GB.
4.6 Inference
Figure 11 presents the inference throughput and memory consumption of models compressed from Llama-7B on a single A100 GPU, using GSM8K as the evaluation set. The results show that low-rank compression methods, including SVD, FWSVD, and Basel, lead to reduced memory consumption and improved throughput as the model size decreases. Throughput and memory usage are primarily dependent on model size, with no significant differences between the methods at equivalent sizes. However, since our proposed Basel method achieves a greater reduction in model size while maintaining similar accuracy to SVD and FWSVD, it improves throughput by up to 19% and reduces memory consumption by up to 37%.
Рис. 11: Throughput and memory consumption of compressed models.
@@ 4.7 Ablation Study
[…]
(a) GSM8k
(b) MATH
Рис. 12: Ablation study: Impact of varying the additional dimension in Basel on the compression of Llama 2-7B for the mathematical reasoning task.
(a) GSM8k
(b) MATH
Рис. 13: Ablation study: Impact of varying the pruning times in Basel on the compression of Llama 2-7B for the mathematical reasoning task.
(a) GSM8K
(b) MATH
Рис. 14: Ablation study: Pass@1 accuracy and size of Llama 2-7B compressed by Basel (proposed) and Basel (free basis) on mathematical reasoning.
Similarly, performing light pruning multiple times allows the model to gradually adapt to parameter reduction, which is especially helpful for extreme compression. Figure 13 illustrates this effect by comparing two pruning strategies—100 times versus 2. The model pruned 100 times consistently achieves higher accuracy at compression ratios greater than 4.
Basel freezes the basis vectors and updates only the singular values (and any additional bases) during compression. To evaluate the benefit of this design choice, we compare it with a variant of Basel where the basis vectors are also updated. Figure 14 presents the performance of both approaches on compressing Llama 2-7B for mathematical reasoning. Freezing the bases results in better accuracy. It also reduces compression time by 33%, as backpropagation involves fewer trainable parameters.
4.8 Compression vs. Training from Scratch
In many real-world LLM deployment scenarios, compute and hardware constraints impose a target model size tailored to a specific application, which may not match any existing pretrained models. For example, Llama-2 offers 7B, 13B, and 70B variants, but no 4B option. To obtain a model of the required size for the target application, one can either (1) train a new model from scratch or (2) compress a nearby larger model. The first approach is often impractical due to high computational costs—training Llama reportedly requires up to one million A100 GPU hours Touvron et al. (2023a)—and limited access to pretraining data. In contrast, model compression is far more efficient; low-rank methods such as Basel typically require less than 0.01% of the pretraining compute.
To assess the effectiveness of compression, we compare a fine-tuned Llama 3.2-1B model with a compressed version of Llama 3.2-3B reduced to 1B parameters using Basel. As shown in Table 2, the compressed model achieves similar performance on mathematical reasoning. This indicates that, in this setting, compression can serve as a viable alternative to training from scratch—offering substantial savings in compute and eliminating the need for pretraining data. These results demonstrate the potential of Basel to support efficient model scaling under deployment constraints.
Таблица 2: Pass@1 accuracy and model size of Llama 3.2-3B compression via Basel for mathematical reasoning.
3B 1B Basel-compressed-3B
Model Size (billion) 3.21 1.24 1.24
GSM8K Acc (%) 72.5 54.4 55.3
MATH Acc (%) 26.1 17.6 16.7
@@ 5 Conclusion
The significant size of large language models leads to high inference costs and demands substantial computing resources. To mitigate these issues, we focus on compressing large language models to meet the specific requirements of target applications. Our approach involves examining these models through the lens of matrix factorization. By viewing the weight matrix of large language models as a linear combination of a group of bases, we have identified that pretrained models often contain many redundant bases that are less useful for target applications. To address this, we propose Basel, a compression algorithm that evaluates the importance of each base for target applications and prunes those that are less significant. Experimental results demonstrate that Basel significantly outperforms state-of-the-art low-rank compression algorithms in achieving deep compression. Basel greatly reduces the inference cost of large language models, making them more accessible and practical for a wider range of applications. This advancement has the potential to democratize the use of large language models, facilitating their adoption and integration across diverse fields and industries.
Список литературы
  • Acharya et al. (2019) Anish Acharya, Rahul Goel, Angeliki Metallinou, and Inderjit Dhillon. Online Embedding Compression for Text Classification Using Low Rank Matrix Factorization. In AAAI, 2019.
  • An et al. (2024) Yongqi An, Xu Zhao, Tao Yu, Ming Tang, and Jinqiao Wang. Fluctuation-based Adaptive Structured Pruning for Large Language Models. In AAAI, 2024.
  • Austin et al. (2021) Jacob Austin, Augustus Odena, Maxwell Nye, Maarten Bosma, Henryk Michalewski, David Dohan, Ellen Jiang, Carrie Cai, Michael Terry, Quoc Le, and Charles Sutton. Program Synthesis with Large Language Models. arXiv preprint arXiv:2108.07732, 2021.
  • Chen et al. (2021a) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, et al. Evaluating Large Language Models Trained on Code. arXiv preprint arXiv:2107.03374, 2021a.
  • Chen et al. (2018) Patrick Chen, Si Si, Yang Li, Ciprian Chelba, and Cho-Jui Hsieh. GroupReduce: Block-Wise Low-Rank Approximation for Neural Language Model Shrinking. In NeurIPS, 2018.
  • Chen et al. (2021b) Patrick Chen, Hsiang-Fu Yu, Inderjit Dhillon, and Cho-Jui Hsieh. DRONE: Data-Aware Low-Rank Compression for Large NLP Models. In NeurIPS, 2021b.
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. Training Verifiers to Solve Math Word Problems. arXiv preprint arXiv:2110.14168, 2021.
  • Denton et al. (2014) Emily L Denton, Wojciech Zaremba, Joan Bruna, Yann LeCun, and Rob Fergus. Exploiting Linear Structure within Convolutional Networks for Efficient Evaluation. NeurIPS, 2014.
  • Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. QLoRA: Efficient Finetuning of Quantized LLMs. In NeurIPS, 2023.
  • Golub & Van Loan (1996) Gene H. Golub and Charles F. Van Loan. Matrix Computations. Johns Hopkins University Press, Baltimore and London, 1996.
  • Hendrycks et al. (2021) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt. Measuring Mathematical Problem Solving with the MATH Dataset. NeurIPS, 2021.
  • Hsu et al. (2022) Yen-Chang Hsu, Ting Hua, Sungen Chang, Qian Lou, Yilin Shen, and Hongxia Jin. Language Model Compression with Weighted Low-Rank Factorization. In ICLR, 2022.
  • Jaderberg et al. (2014) Max Jaderberg, Andrea Vedaldi, and Andrew Zisserman. Speeding Up Convolutional Neural Networks with Low Rank Expansions. arXiv preprint arXiv:1405.3866, 2014.
  • Noach & Goldberg (2020) Matan Ben Noach and Yoav Goldberg. Compressing pre-trained language models by matrix decomposition. In AACL/IJCNLP, 2020.
  • Povey et al. (2018) Daniel Povey, Gaofeng Cheng, Yiming Wang, Ke Li, Hainan Xu, Mahsa Yarmohammadi, and Sanjeev Khudanpur. Semi-Orthogonal Low-Rank Matrix Factorization for Deep Neural Networks. In INTERSPEECH, 2018.
  • Sharma et al. (2024) Pratyusha Sharma, Jordan T. Ash, and Dipendra Misra. The Truth is in There: Improving Reasoning in Language Models with Layer-Selective Rank Reduction. In ICLR, 2024.
  • Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. Llama: Open and Efficient Foundation Language Models. arXiv preprint arXiv:2302.13971, 2023a.
  • Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. Llama 2: Open Foundation and Finetuned Chat Models. arXiv preprint arXiv:2307.09288, 2023b.
  • Xue et al. (2013) Jian Xue, Jinyu Li, and Yifan Gong. Restructuring of Deep Neural Network Acoustic Models with Singular Value Decomposition. In INTERSPEECH, 2013.
  • Yu & Wu (2023) Hao Yu and Jianxin Wu. Compressing Transformers: Features Are Low-Rank, But Weights Are Not! In AAAI, 2023.
  • Zhang et al. (2015) Xiangyu Zhang, Jianhua Zou, Kaiming He, and Jian Sun. Accelerating very deep convolutional networks for classification and detection. IEEE Transactions on Pattern Analysis and Machine Intelligence, 38(10):1943–1955, 2015.